• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name S. Sirca

    Log entry time 00:12:53 on January20,2001

    Entry number 55206

    keyword=Evening shift summary


    Run# Start/Stop Events Comments
    1519 16:14/16:26 0.38 M OK, LD2, radiator OUT, beam off for accel re-tune
    16:38 Beam is back, take ~44muA beam
    1520 16:39/17:04 1.00 M OK
    1521 17:04/17:30 1.00 M OK
    1522 17:30/18:00 1.00 M OK
    18:00 Replace LD2 by LH2, adjust JT valves
    1523 18:26/18:49 1.00 M OK, LH2, 6% radiator IN, ~48muA beam
    1524 18:50/19:10 1.00 M OK
    1525 19:10/19:32 1.00 M OK
    1526 19:35/20:23 1.00 M OK, LH2, radiator OUT, ~48muA beam
    1527 20:24/20:36 0.25 M OK
    1528 20:52/21:09 0.14 M OK, 15 cm dummy target with 30 muA
    21:10/22:10 Switching to kinematics #17, LH2, radiator OUT
    1529 22:10/22:22 junk, realised left scint HVs were off
    1530 22:22/22:27 short run, beam off after raster IOC crash
    22:27 MCC makes hall access to recover IOC
    23:59 Still no time estimate for restored beam